Juwan is 49 years old. Katie is 16 years old. How many years until Juwan's age is 2 times Katie's age?
Answer by ikleyn(52787)

.
49 + x = 2*(16+x)


49 + x = 32 + 2x


49 - 32 = 2x - x


x = 17.


Answer.  17 years.